OpenStudy (anonymous):

ASAP: Please help!? The area of this cylinder's base is one hundred fifty square inches. The height of the cylinder is 6 inches. What is the volume of the cone that has the same base area and height? Show the steps of your solution and explain your work. cylinder having a height of 6 inches

OpenStudy (anonymous):

what is the formlula for volume for cone? 1/3 pi r^2 h

OpenStudy (anonymous):

pi r^2 part we know it to be 50 h= 6

OpenStudy (anonymous):

1/3

OpenStudy (nurali):

The formula for the volume of a cylinder is: v1 = π(r^2)h and πr^2 is the same thing as the base area. Since we're given that that is 150, we can find the radius: πr^2 = 150 r^2 = 150/π r = √(150/π) And we're given that the height is 6 The volume of a cone is: V = (π/3)(r^2)h Plugging in r = √(150/π) and h = 6 V = (π/3)(√(150/π)^2)(6) V = (π/3)(150/π)(6) V = 300 300 cubic inches

OpenStudy (anonymous):

Thank you ♥

OpenStudy (nurali):

My Pleasure.